# From Freshness to Effectiveness: Goal-Oriented Sampling for Remote Decision Making

### v2

@@ III System Model and Problem Formulation
We consider a time-slotted age-aware remote MDP problem illustrated in Fig. 1(c). Let $X_{t}\in\mathcal{S}$ be the controlled source of interest at time slot $t$. The evolution of the source is a Markov decision process, characterized by the transition probability $\Pr(X_{t+1}|X_{t},a_{t})$ [^3] For short-hand notations, we use the transition probability matrix $\mathbf{P}_{a}$ to encapsulate the dynamics of the source given an action $a_{t}=a$., where $a_{t}\in\mathcal{A}$ represents the controlled action taken by the remote decision maker to control the source in the desired way. The sampler conducts the sampling action $u_{t}\in\{0,1\}$, with $u_{t}=1$ representing the sampling action and $u_{t}=0$ otherwise. Let $S_{i}$ be the sampling time of the $i$-th delivered packet, and $D_{i}$ be the corresponding delivery slot. The random channel delay of the $i$-th packet is denoted as $Y_{i}\in\mathcal{Y}\subseteq\mathbb{N}^{+}$, which is independent of the source $X_{t}$ and is bounded $\max[{Y_{i}}]<\infty$. The sampling times $S_{0},S_{1},\cdots$ shown in Fig. 2 record the time stamp when $u_{t}=1$, given by
$S_{i}=\max\{t\left|t\leq D_{i},u_{t}=1\right.\},\forall i\in\mathbb{N},$ (2)
@@
At the sampling time $S_{i},\forall i\in\mathbb{N}$, the state $X_{S_{i}}$ along with the corresponding time stamp $S_{i}$ is encapsulated into a packet $(X_{S_{i}},S_{i})$, which is transmitted to a remote decision maker. Upon the receipt of the packet $(X_{S_{i}},S_{i})$ during $t\in[D_{i},D_{i+1})$, the observation history at the decision maker is $\{(X_{S_{j}},S_{j}):{j\leq i}\}$. By employing (1), this sequence can be equivalently represented as $\{(X_{k-\Delta(k)},\Delta(k)):{k\leq t}\}$.
Let $\mathcal{I}_{t}=\{(X_{k-\Delta(k)},\Delta(k),a_{k-1}^{S},a_{k-1}):{k\leq t}\}$ denote the observation history and action history (or simply history) available to the decision maker up to time $t$. The decision maker is tasked with determining both the sampling action and the controlled action $(u_{t},a_{t})\in\{0,1\}\times\mathcal{A}$ at each time slot $t$ by leveraging the history $\mathcal{I}_{t}$. A decision policy of the decision maker is defined as a mapping from the history to a distribution over the joint action space $\{0,1\}\times\mathcal{A}$, denoted by
$\pi_{t}:\mathcal{I}_{t}\rightarrow\mathcal{P}(\mathcal{A}\times\{0,1\}),$ (3)
where $\mathcal{P}(\cdot)$ is a simplex space which represents the probability that an action is taken. Similar to [63], we assume that the sampling policy satisfies the following two conditions:
  • $i$) No sample is taken when the channel is busy, , $S_{i+1}\geq D_{i}$, or equivalently,
    $S_{i+1}-Z_{i}=D_{i}\text{ with }Z_{i}\geq 0\text{ and }i\in\mathbb{N},$ (4)
    with $Z_{i}$ representing the sampling waiting time. From this assumption, we can also obtain that $D_{i}=S_{i}+Y_{i},\forall i\in\mathbb{N}$.
@@
  • $ii$) The inter-sample times $G_{i}=S_{i+1}-S_{i}$ is a regenerative process [87, Section 6.1]: There is a sequence $0\leq g_{1}<g_{2}<\cdots$ of almost surely finite random integers such that the post-$g_{j}$ process $\{G_{{g_{j}}+i}\}_{i\in\mathbb{N}}$ has the same distribution as the post-$g_{1}$ process $\{G_{{g_{1}}+i}\}_{i\in\mathbb{N}}$ and is independent of the pre-$g_{j}$ process $\{G_{i}\}_{i\in\{0,1,\cdots g_{j}-1\}}$. Condition $ii$) implies that, almost surely [^4] This assumption also implies that the waiting time $Z_{i}$ is bounded, belonging to a subset of nature numbers with $Z_{i}\in\mathcal{Z}\subseteq\mathbb{N}$.
    $\lim\limits_{i\rightarrow\infty}S_{i}=\infty,~{}~{}~{}\lim\limits_{i\rightarrow\infty}D_{i}=\infty.$ (5)
In addition, we assume that the controlled policy satisfies the following condition: the action $a_{t}$ is updated only upon the delivery of a sample $X_{S_{i}}$ [^5] Throughout the time interval $t\in[D_{i},D_{i+1})$, the decision maker’s observations are fixed and consist only of $\{(X_{S_{j}},S_{j}):{j\leq i}\}$. Thus, we assume that the chosen actions remain constant in this period. , i.e.,
$a_{t}=A_{i},D_{i}\leq t\leq D_{i+1},i\in\mathbb{N},$ (6)
where $A_{i}\in\mathcal{A}$ is the updated controlled action upon the delivery of packet $(X_{S_{i}},S_{i})$. We consider a bounded cost function $\mathcal{C}(X_{t},a_{t})<\infty$, which represents the immediate cost incurred when action $a_{t}$ is taken in state $X_{t}$.
@@
Under the above assumptions, the objective of the system is to design the optimal joint sampling and decision policies at each time slot, i.e., $\pi_{0},\pi_{1},\pi_{2}\cdots$, to minimize the long-term average cost, subject to a long-term average sampling frequency constraint:
@@
[…]
@@
We remark that we will study Problem 1 both with (see Section IV) and without (see Section VI) the sampling frequency constraint (7b). To distinguish these two problems, we use $h^{\star}$ to denote the optimal value of the problem with the rate constraint and $\rho^{\star}$ to denote the optimal value of the problem without the sampling frequency constraint.
In this section, we address the unconstrained problem to determine $\rho^{\star}$. A series of theoretical results are developed through a divide-and-conquer approach. In subsection IV-A, we identify that the input to the policy $\pi_{t}$ is subject to the curse of history [88]. To address this, we analyze the sufficient statistics of the input to extract the most compressed representation that is necessary for decisions. In subsection IV-B, we rewrite the unconstrained problem into a non-linear fractional program. By utilizing a Dinkelbach-like method [89], we transform the non-linear fractional programming into an infinite horizon MDP problem given a Dinkelbach parameter. The problem then becomes finding the Dinkelbach parameter such that the optimal value of the standard MDP is zero, which is equivalently a root-finding problem. To search for the root, in subsection IV-C we review a typical two-layer nested algorithm, namely Bisec-RVI [1, Algorithm 2] [^6] The idea of the two-layer nested algorithm has been applied in [63, 20, 90] and [44] to achieve Age-optimal and Mean Square Error (MSE)-optimal sampling and scheduling.. We note that the inner-layer Relative Value Iteration (RVI) algorithm of Bisec-RVI may suffer from divergence and thus propose an improved Bisec-$\tau$-RVI algorithm to achieve provable convergence.
IV-A Sufficient Statistics of History
The policy $\pi_{t}$ maps the history $\mathcal{I}_{t}$ to the action simplex space $\mathcal{P}(\mathcal{A}\times\{0,1\})$. A significant challenge in solving $\pi_{t}$ is that the size of the history space $\mathcal{I}_{t}=\{(X_{k-\Delta(k)},\Delta(k),a_{k-1}^{S},a_{k-1}):{k\leq t}\}$ explodes exponentially as $t$ increases, known as the curse of history [88]. In this subsection, we analyze the sufficient statistic of $\mathcal{I}_{t}$ to address the curse of history. In decision theory, a sufficient statistic is defined as:
@@ IV-A Sufficient Statistics of History
This definition suggests that the decision making leveraging the sufficient statistics $\mathcal{H}_{t}(\mathcal{I}_{t})$ can achieve the same performance as using the complete history $\mathcal{I}_{t}$. Thus, the compression $\mathcal{H}_{t}(\mathcal{I}_{t})$ is sufficient for the agent to implement an optimal action, enabling the design of efficient policies that maintain optimality while overcoming the curse of history. The following Lemma establishes an efficient sufficient statistics of the history $\mathcal{I}_{t}$ in a piecewise manner.
@@
The proof is detailed in the the short version of this work in [1, Appendix A]. ∎
The above Lemma 1 indicates that determining the optimal policy
$\pi_{t}:\mathcal{I}_{t}\rightarrow\mathcal{P}(\mathcal{A}\times\{0,1\})$ (9)
for Problem 1 is equivalent to solving for an alternative policy given as
$\phi_{t}:\mathcal{S}\times\mathcal{Y}\times\mathcal{A}\rightarrow\mathcal{P}(\mathcal{A}\times\mathcal{Z}),\forall t\in[D_{i},D_{i+1}),i\in\mathbb{N},$ (10)
which is a mapping from the sufficient statistics $\mathcal{G}_{i}=(X_{S_{i}},Y_{i},A_{i-1})\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ to a distribution over the joint sampling and decision-making actions $(A_{i},Z_{i})\in\mathcal{A}\times\mathcal{Z}$. Because the Cartesian product $\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ is independent of time $t$, the exponential growth of $\mathcal{I}_{t}$ with respect to $t$ is thus mitigated. Consequently, the policy $\{\pi_{t}\}_{t=0}^{\infty}$ in Problem 1 can be replaced by an alternative policy $\{\phi_{t}\}_{t=0}^{\infty}$, which reformulates the unconstrained version of Problem 1 as:
@@
$\rho^{\star}\triangleq\inf_{\boldsymbol{\psi}}\limsup_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\mathcal{C}(X_{t},a_{t})\right].$ (11)
Here, we use $\boldsymbol{\psi}$ to denote the policy $\{\phi_{t}\}_{t=0}^{\infty}$.
IV-B A Reformulation of Problem 2
As indicated in Lemma 1, the sufficient statistic of history is in a piecewise manner with respect to the time slot $t$. This allows us to shift our problem-solving perspective from a slot-by-slot analysis to an interval-by-interval approach. Assuming that the inter-sampling times $G_{i}$ constitute a regenerative process and given $\lim\limits_{i\rightarrow\infty}D_{i}=\infty$, we can reformulate the objective function in Problem 2 as:
$\limsup_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\mathcal{C}(X_{t},a_{t})\right]$ (12a)
$=\lim_{\it{D}_{\mathrm{n}}\rightarrow\infty}\frac{\mathbb{E}\left[\sum_{t=1}^{\it{D}_{\mathrm{n}}}\mathcal{C}(X_{t},a(t))\right]}{\mathbb{E}\left[\it{D}_{\mathrm{n}}\right]}$ (12b)
$=\lim_{\mathrm{n}\rightarrow\infty}\frac{\sum_{i=0}^{n-1}\mathbb{E}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]}{\sum_{i=0}^{n-1}\mathbb{E}\left[\it{D}_{i+1}-\it{D}_{i}\right]}$ (12c)
$=\lim_{\mathrm{n}\rightarrow\infty}\frac{\sum_{i=0}^{n-1}\mathbb{E}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]}{\sum_{i=0}^{n-1}\mathbb{E}\left[Y_{i+1}+Z_{i}\right]}.$ (12d)
We can then rewrite Problem 2 as the following interval-by-interval variant:
Problem 3 (Interval-by-Interval Reformulation).
$\rho^{\star}\triangleq\inf_{\boldsymbol{\psi}}\lim_{\mathrm{n}\rightarrow\infty}\frac{\sum_{i=0}^{n-1}\mathbb{E}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]}{\sum_{i=0}^{n-1}\mathbb{E}\left[Y_{i+1}+Z_{i}\right]}.$ (13)

### v3

@@ III System Model and Problem Formulation
We consider a time-slotted [^4] As the proposed AR-MDP is formulated as an extension of the discrete-time MDP framework, a time-slotted system model is employed to maintain structural consistency with MDP. Therefore, all key variables including transmission delay $Y_{i}$, AoI $\Delta(k)$, and sampling time $S_{i}$ are accordingly defined and evolve over discrete time slots. age-aware remote MDP problem illustrated in Fig. 1(c). Let $X_{t}\in\mathcal{S}$ be the controlled source of interest at time slot $t$. The evolution of the source is a Markov decision process, characterized by the transition probability $\Pr(X_{t+1}|X_{t},a_{t})$ [^5] For short-hand notations, we use the transition probability matrix $\mathbf{P}_{a}$ to encapsulate the dynamics of the source given an action $a_{t}=a$., where $a_{t}\in\mathcal{A}$ represents the controlled action taken by the remote decision maker to control the source in the desired way. We assume that both the state space $\mathcal{S}$ and the action space $\mathcal{A}$ are finite. This finite setting is a common starting point in infinite-horizon average-cost MDP literature [76, 12] and allows the proposed AR-MDP to inherit well-established optimality and convergence results. We view AR-MDP as a stepping stone toward more general models, and refer readers to [12, Chap. 4.6] for approaches that can be used to extend our framework to infinite spaces.
The sampler conducts the sampling action $u_{t}\in\{0,1\}$, with $u_{t}=1$ representing the sampling action and $u_{t}=0$ otherwise. Let $S_{i}$ be the sampling time of the $i$-th delivered packet, and $D_{i}$ be the corresponding delivery slot. The random channel delay of the $i$-th packet is denoted as $Y_{i}\in\mathcal{Y}\subseteq\mathbb{N}^{+}$, which is independent of the source $X_{t}$ and is bounded $\max[{Y_{i}}]<\infty$. The sampling times $S_{0},S_{1},\cdots$ record the time stamp when $u_{t}=1$, given by
$S_{i}=\max\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t\in\mathbb{N}}\left|t\leq D_{i},u_{t}=1\right.\},\quad\forall i\in\mathbb{N},$ (3)
@@
Since the system is time-slotted, the AoI evolves in discrete steps. Specifically, for any slot $t$, let $i(t)\triangleq\max\{i:D_{i}\leq t\}$ be the index of the most recently delivered update. Then $\Delta(t)=t-S_{i(t)}$. Hence, upon a delivery at $t=D_{i}$, the AoI drops to
$\Delta(D_{i})=D_{i}-S_{i}=Y_{i},$ (4)
which is not necessarily zero unless the channel delay is zero. Moreover, during the interval $t\in[D_{i},D_{i+1})$, no new observation is delivered and the age $\Delta(t)$ drifts with $\Delta(t+1)=\Delta(t)+1$. The discrete-time AoI dynamics are shown in Fig. 2.
@@
At the sampling time $S_{i},\forall i\in\mathbb{N}$, the state $X_{S_{i}}$ along with the corresponding time stamp $S_{i}$ is encapsulated into a packet $(X_{S_{i}},S_{i})$, which is transmitted to a remote decision maker. Upon receipt of the packet $(X_{S_{i}},S_{i})$ at delivery time slot $D_{i}$, the observation history at the decision maker is $\{(X_{S_{j}},S_{j},D_{j}):j\leq i\},t\in[D_{i},D_{i+1}).$ By employing (1), for any time slot $t\in\mathbb{N}$, the freshest available sample at the decision maker is $(X_{t-\Delta(t)},t-\Delta(t))$. As $t$ is known to the decision maker, the observation history up to time slot $t$ is equivalently expressed by AoI, given as: $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{(X_{k-\Delta(k)},\Delta(k)):k\leq t,k\in\mathbb{N}\}.$
III-A Protocol and Assumptions
Similar to [88], we impose the following assumptions in sampling:
  • (S1)
    A new sample cannot be generated until the previous sample has been delivered. Specifically,
    $S_{i+1}=D_{i}+Z_{i},\quad Z_{i}\geq 0,\quad i\in\mathbb{N},$ (5)
    where $Z_{i}$ is the sampling waiting time after the delivery at $D_{i}$. Consequently, the delivery time satisfies $D_{i}=S_{i}+Y_{i}$ for all $i\in\mathbb{N}$, where $Y_{i}$ is the random delay of the $i$-th sample. Moreover, under (S1) and the delivery timeline, the observation is piecewise constant between two consecutive deliveries. Specifically, for any $t\in[D_{i},D_{i+1})$ we have $t-\Delta(t)=S_{i}$, and hence $O(t)=X_{t-\Delta(t)}=X_{S_{i}},D_{i}\leq t<D_{i+1},$ while the staleness $\Delta(t)$ increases deterministically within the interval.
  • (S2)
@@ III-A Protocol and Assumptions
    The inter-sample times $G_{i}=S_{i+1}-S_{i}$ form a regenerative process [36, Section 6.1]. Hence, almost surely [^6] This assumption also implies that the waiting time $Z_{i}$ is bounded, belonging to a subset of nature numbers with $Z_{i}\in\mathcal{Z}\subseteq\mathbb{N}$.,
    $\lim_{i\to\infty}S_{i}=\infty,\qquad\lim_{i\to\infty}D_{i}=\infty.$ (6)
In addition, we adopt a holding-action paradigm consistent with the timeline in Fig. 2:
  • (A1)
    The controlled action is updated only upon the delivery of a sample. That is, upon the delivery time $D_{i}$, the decision maker selects an updated action $A_{i}\in\mathcal{A}$ based on the available history, and then holds it constant until the next delivery:
    $a_{t}=A_{i},\quad D_{i}\leq t<D_{i+1},\quad i\in\mathbb{N}.$ (7)
    As a result, the underlying state continues to evolve according to the controlled Markov kernel, i.e., $X_{t+1}\sim P(\cdot\mid X_{t},A_{i})$ for all $t\in[D_{i},D_{i+1})$.
Remark 1.
The modeling assumption (A1) is widely used in networked and remote control systems where the controller can revise its command only when a new measurement is received; see, e.g., [47, 27, 38]. Typical real-world systems that follow the holding-action paradigm include:
  • Robotic manipulation and teleoperation: In robotic teleoperation under constant communication delays, the controller continuously updates the motion command based on the current local state and the most recently received delayed remote state, effectively maintaining past remote information during delay periods [47].
  • Multi-UAV or vehicular coordination: each autonomous agent maintains its last chosen coordination strategy (e.g., formation control gain or following distance target) until updated state information is received from neighboring agents [27].
  • Industrial supervisory control: in process plants or smart grids, a supervisory controller holds the previously assigned operation mode (e.g., heater on/off state, pump flow rate setpoint) during communication gaps between control center and local devices [38].
Remark 2 (Beyond holding actions).
The holding-action rule (7) induces an epoch-based decision structure: between two consecutive deliveries, no new observation is received and the applied action is fixed, and the state continues to evolve according to the controlled Markov kernel under the held action. This structure is essential for compressing the growing history into a finite-dimensional sufficient statistic (Lemma 1) and for constructing a finite-state lifted MDP amenable to our average-cost optimality analysis and algorithms [^7] If (A1) is relaxed so that the decision maker can adapt $a_{t}$ within $t\in[D_{i},D_{i+1})$ based on the deterministic AoI drift, then the sufficient statistic generally becomes belief-based: the conditional distribution of $X_{t}$ depends on the staleness and the within-epoch action sequence. This leads to a (belief-)MDP/POMDP formulation with either a continuous state space (belief simplex) or a significantly enlarged epoch action space (age-indexed action plans). We leave this non-holding extension as future work..
III-B Joint Sampling and Decision-Making Policy
Let $\mathcal{I}_{t}=\{(X_{k-\Delta(k)},\Delta(k),u_{k-1},a_{k-1}):k\leq t\}$ denote the history available to the decision maker up to time $t$. Under the protocol in (5) and (7), the slot-level actions $(u_{t},a_{t})$ are induced by the epoch-level decisions $\{(A_{i},Z_{i})\}$ made at delivery epochs. A possibly randomized decision policy is a sequence of mappings from the history to a distribution over the joint action space $\{0,1\}\times\mathcal{A}$:
$\pi_{t}:\mathcal{I}_{t}\rightarrow\mathcal{P}(\{0,1\}\times\mathcal{A}),$ (8)
where $\mathcal{P}(\cdot)$ is a simplex space which represents the probability that an action is taken.
An epoch-based policy is defined as
$\phi:\mathcal{S}\times\mathcal{Y}\times\mathcal{A}\to\mathcal{P}(\mathcal{A}\times\mathcal{Z}),$ (9)
which induces the slot-level policy $\pi_{t}$ via $S_{i+1}=D_{i}+Z_{i}$ and $a_{t}=A_{i}$ on $D_{i},D_{i+1})$. For completeness, a slot-level policy can be written as $\pi_{t}:\mathcal{I}_{t}\to\mathcal{P}(\{0,1\}\times\mathcal{A})$, which in our setting is induced by $\phi$. (See Lemma 1).
@@ III-B Joint Sampling and Decision-Making Policy
We consider a bounded cost function $\mathcal{C}(X_{t},a_{t})<\infty$, which represents the immediate cost incurred when action $a_{t}$ is taken in state $X_{t}$. Under the above assumptions, the objective of the system is to design the optimal joint sampling and decision policies at each time slot, i.e., $\pi_{0},\pi_{1},\pi_{2}\cdots$, to minimize the long-term average cost, subject to a long-term average sampling frequency constraint:
@@
$\text{s.t. }\liminf_{N\to\infty}\frac{1}{N}\mathbb{E}_{\pi_{0:\infty}}\left[\sum_{i=0}^{N-1}(S_{i+1}-S_{i})\right]\geq\frac{1}{f_{\text{max}}},$ (10b)
@@
We remark that we will study Problem 1 both without (see Section IV and V) and with (see Section VI and VII) the sampling frequency constraint. To distinguish these two problems, we use $h^{\star}$ to denote the optimal value of the problem with the rate constraint and $\rho^{\star}$ to denote the optimal value of the problem without the sampling frequency constraint.
III-C Sufficient Statistics of History
In principle, the joint policy $\pi_{t}$ maps the growing history $\mathcal{I}_{t}$ to a distribution over $(u_{t},a_{t})$, which leads to the classical curse of history and makes direct dynamic programming intractable.
More importantly, under delayed observations, optimal decisions generally depend not only on the latest delivered sample value but also on its staleness. In our time-slotted model, the staleness at delivery epochs is $\Delta(D_{i})=Y_{i}$. Within the interval $t\in[D_{i},D_{i+1})$, the freshest delivered sample remains $X_{S_{i}}$ and the AoI drifts deterministically as $\Delta(t)=Y_{i}+(t-D_{i})$, i.e., the time since sampling equals $t-S_{i}=\Delta(t)$. This determines how many Markov transitions have occurred since the sampled state. Consequently, even conditioned on the same delivered sample $X_{S_{i}}$ and held actions, different $Y_{i}$ (and hence different $\Delta(t)$ in the epoch $t\in[D_{i},D_{i+1})$) induce different conditional distributions (beliefs) of the current state.
Crucially, under the holding-action rule (7), no new observation is delivered between two consecutive deliveries and the action is kept constant. This induces an epoch-based decision structure, under which the growing history can be compressed into a finite-dimensional sufficient statistic. The resulting statistic, given in Lemma 1, serves as the state of the lifted MDP and enables our subsequent analysis and algorithms.
A sufficient statistic is defined as follows.
@@ III-C Sufficient Statistics of History
This definition suggests that decision-making that leverages the sufficient statistics $\mathcal{H}_{t}(\mathcal{I}_{t})$ can achieve the same performance as using the complete history $\mathcal{I}_{t}$. Thus, the compression $\mathcal{H}_{t}(\mathcal{I}_{t})$ is sufficient for the agent to implement an optimal action, enabling the design of efficient policies that maintain optimality while overcoming the curse of history. In particular, the time stamp carried by each packet makes the staleness observable at delivery: upon receiving $(X_{S_{i}},S_{i})$ at time $D_{i}$, the decision maker can compute $\Delta(D_{i})=D_{i}-S_{i}=Y_{i}$. This staleness information enters the sufficient statistic and therefore the optimal delivery-epoch policy. The following Lemma establishes efficient sufficient statistics of the history $\mathcal{I}_{t}$ in a piecewise manner.
@@
See Appendix A.
Lemma 1 indicates that solving Problem 1 over the history-dependent slot-level policies
$\pi_{t}:\mathcal{I}_{t}\rightarrow\mathcal{P}(\{0,1\}\times\mathcal{A})$ (12)
is equivalent to solving over delivery-epoch policies that depend only on $\mathcal{G}_{i}$:
$\phi_{i}:\mathcal{S}\times\mathcal{Y}\times\mathcal{A}\rightarrow\mathcal{P}(\mathcal{A}\times\mathcal{Z}),\quad\forall i\in\mathbb{N},$ (13)
which maps the sufficient statistic $\mathcal{G}_{i}=(X_{S_{i}},Y_{i},A_{i-1})$ to a distribution over the joint epoch actions $(A_{i},Z_{i})\in\mathcal{A}\times\mathcal{Z}$. Consequently, the AoI at delivery, equivalently the realized delay $Y_{i}=\Delta(D_{i})$, is an explicit argument of the optimal policy via $\mathcal{G}_{i}$. This provides a concrete AoI-related conclusion: even for the same delivered sample value $X_{S_{i}}$, different staleness values $Y_{i}$ can induce different beliefs and thus the optimal epoch sampling and decision-making may depend on $Y_{i}$. Because the Cartesian product $\mathcal{S}\times\mathcal{Y}\times\mathcal{A}$ does not grow with time, this reformulation avoids the exponential growth of the original information set $\mathcal{I}_{t}$. From assumptions (S1) and (S2), we reformulate Problem 1 as:
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\inf_{\boldsymbol{\psi}}\limsup_{N\rightarrow\infty}\frac{\mathbb{E}_{\boldsymbol{\psi}}\left[\sum_{i=0}^{N-1}\sum_{t=D_{i}}^{D_{i+1-1}}\mathcal{C}(X_{t},a_{t})\right]}{\mathbb{E}_{\boldsymbol{\psi}}[D_{N}]}}$ (14a)
$\text{s.t. }\liminf_{N\to\infty}\frac{1}{N}\mathbb{E}_{\boldsymbol{\psi}}\left[\sum_{i=1}^{N}(S_{i+1}-S_{i})\right]\geq\frac{1}{f_{\text{max}}},$ (14b)
Here, we use $\boldsymbol{\psi}$ to denote the policy $\{\phi_{i}\}_{i=0}^{\infty}$.
In this section, we address the unconstrained problem to determine $\rho^{\star}$. A series of theoretical results are developed through a divide-and-conquer approach. In subsection IV-A, we rewrite the unconstrained problem into a non-linear fractional program. By utilizing a Dinkelbach-like method [28], we transform the non-linear fractional programming into an infinite horizon MDP problem given a Dinkelbach parameter. The problem then becomes finding the Dinkelbach parameter such that the optimal value of the standard MDP is zero, which is equivalently a root-finding problem. To search for the root, in subsection IV-B we review a typical two-layer nested algorithm, namely Bisec-RVI [49, Algorithm 2] [^8] The idea of the two-layer nested algorithm has been applied in [88, 85, 9] and [84] to achieve Age-optimal and Mean Square Error (MSE)-optimal sampling and scheduling.. We note that the inner-layer Relative Value Iteration (RVI) algorithm of Bisec-RVI may suffer from divergence and thus propose an improved Bisec-$\tau$-RVI algorithm to achieve provable convergence.
IV-A A Reformulation of Problem 2
From the action-holding assumption (A1), we can rewrite the objective function in Problem 2 as:
$\limsup_{N\rightarrow\infty}\frac{\mathbb{E}_{\boldsymbol{\psi}}\left[\sum_{i=0}^{N-1}\sum_{t=D_{i}}^{D_{i+1-1}}\mathcal{C}(X_{t},a_{t})\right]}{\mathbb{E}_{\boldsymbol{\psi}}[D_{N}]}$ (15a)
$=\lim_{\mathrm{n}\rightarrow\infty}\frac{\sum_{i=0}^{N-1}\mathbb{E}_{\boldsymbol{\psi}}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]}{\sum_{i=0}^{N-1}\mathbb{E}_{\boldsymbol{\psi}}\left[\it{D}_{i+1}-\it{D}_{i}\right]}$ (15b)
$=\lim_{\mathrm{n}\rightarrow\infty}\frac{\sum_{i=0}^{n-1}\mathbb{E}_{\boldsymbol{\psi}}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]}{\sum_{i=0}^{n-1}\mathbb{E}_{\boldsymbol{\psi}}\left[Y_{i+1}+Z_{i}\right]}.$ (15c)
We can then express the unconstrained version of Problem 2 as the epoch-by-epoch variant:
Problem 3 (Epoch-by-Epoch Reformulation).
$\rho^{\star}\triangleq\inf_{\boldsymbol{\psi}}\lim_{\mathrm{N}\rightarrow\infty}\frac{\sum_{i=0}^{N-1}\mathbb{E}_{\boldsymbol{\psi}}\left[\sum_{t=\it{D}_{i}}^{\it{D}_{i+1}-1}\mathcal{C}(X_{t},A_{i})\right]}{\sum_{i=0}^{N-1}\mathbb{E}_{\boldsymbol{\psi}}\left[Y_{i+1}+Z_{i}\right]}.$ (16)